Helium gas at 27 ºC and pressure of 1.6 atm is compressed adiabatically to one fourth of its original volume. Find the final pressure (in atm) of the gas. (2sgf)
γ=1.4
1
Expert's answer
2020-11-05T03:53:51-0500

Adiabatic process can be represented by the polytropic process equation:


"p_1V_1^{\gamma} =p_2V_2^{\gamma}"

where "p_1 = 1.6 \space atm" is the initial pressure, "p_2" is the final pressure and "V_1, V_2" are intial and final volumes respectively. They are related to each other in the following way, since the gas was compressed to one fourth of its original volume:


"\dfrac{V_1}{V_2 } = 4"


Expressing the final temperature, obtain:


"p_2 = p_1 \left( \dfrac{V_1}{V_2} \right)^{\gamma} = 1.6\cdot 4 ^{1.4} \approx 11.14 \space atm"

Answer. 11.14 atm.